Grade 12Physical Chemistry

The lattice energy of solid NaCl is 180 kcal/mol. The dissolution of the solid in water in the form of ions is endothermic to the extent of 1 kcal/mol. If the solution energies of Na+ and Cl- are in the ratio 6:5, what is the enthalpy of hydration of Na+ ion.
(A) – 85.6 kcal/mol
(B) – 97.5 kcal/mol
(C) 82.6 kcal/mol
(D) +100 kcal/mol

Shibashis Mallik
11 Years agoGrade 12

1 Answer

Naveen Kumar
11 Years ago
The enthalpy of hydration + energy supplied= enthalpy of lattice
So enthalpy of hydration= 180-1=179 KCal/mol
Let 6y is the energy of hydration of the Na+
So, 6y + 5y=179
y=16.3
So, 6y ~ 97.5 KCal/mol
The sign will be -ve as this energy is released during the process.